### v1

@@ Abstract
This paper provides a rigorous analysis on boundary element methods for the magnetic field integral equation on Lipschitz polyhedra. The magnetic field integral equation is widely used in practical applications to model electromagnetic scattering by a perfectly conducting body. The governing operator is shown to be coercive by means of the electric field integral operator with a purely imaginary wave number. Consequently, the continuous variational problem is uniquely solvable, given that the wave number does not belong to the spectrum of the interior Maxwell’s problem. A Galerkin discretization scheme is then introduced, employing Raviart-Thomas basis functions for the solution space and Buffa-Christiansen functions for the test space. A discrete inf-sup condition is proven, implying the unique solvability of the discrete variational problem. An asymptotically quasi-optimal error estimate for the numerical solutions is established, and the convergence rate of the numerical scheme is examined. In addition, the resulting matrix system is shown to be well-conditioned regardless of the mesh refinement. Some numerical results are presented to support the theoretical analysis.
@@
magnetic field integral equation, MFIE, electromagnetic scattering, boundary element methods, Galerkin boundary element discretization
@@ 1 Introduction
With boundary element methods (BEMs) in mind, we assume that the domain $\Omega$ is a polyhedron with Lipschitz continuous boundary $\Gamma:=\partial\Omega$ that only consists of plane faces $\Gamma_{j}$, $j=1,2,\ldots,N_{F}$. The extension of the results to Lipschitz curvilinear polyhedra is straightforward. On polyhedral boundary $\Gamma$, the definition of Sobolev spaces $\operatorname{H}^{s}(\Gamma)$, in particular with $\left|{s}\right|\geq 1$, and differential operators are not obvious as those on smooth surfaces. By localization to each face $\Gamma_{j}$, Buffa and Ciarlet successfully developed functional setting for Maxwell’s equations on Lipschitz polyhedra, where tangential traces of vector fields in $\operatorname{\mathbf{H}}(\textbf{{curl}},\Omega)$, their trace spaces and relevant surface differential operators were properly characterized [1, 3, 4]. All ingredients necessary to develop a standard functional theory, including integration by parts formulas and Hodge decompositions, were also provided. A generalization for general Lipschitz domains was then presented in [6].
@@
Based on the developed functional theory for Maxwell’s equations on Lipschitz domains, a natural BEM for the EFIE on polyhedra was studied in [22]. The continuous and discrete variational problems are coercive with respect to continuous and semidiscrete Hodge decompositions, respectively. The Galerkin discretization of the EFIE employing Raviart-Thomas boundary elements was analyzed, exhibiting the quasi-optimal convergence of numerical solutions. Another discretization scheme for the EFIE on non-smooth Lipschitz domains was proposed in [5], which is based on standard boundary elements rather than Raviart-Thomas elements. On polyhedral surfaces, an asymptotically quasi-optimal convergence behaviour was also achieved. BEMs for Maxwell transmission problems on Lipschitz domains were investigated in [9]. The Galerkin discretization with Raviart-Thomas spaces was considered, and the results were applied to electromagnetic scattering by perfect electric conductors and by dielectric bodies. In the case of the perfect conductor scattering, only the EFIE was examined, also yielding quasi-optimal numerical solutions. To determine the actual rate of convergence on polyhedral domains, the singularities of solution to Maxwell’s equations must also be taken into account [17]. More specifically, the convergence rate of BEMs for the EFIE is limited by the regularity parameter $s^{\ast}$ of the Laplace-Beltrami operator on $\Gamma$, which depends on the geometry of $\Gamma$ in neighborhoods of vertices only, see, e.g., [22, 5].
@@
In addition to the EFIE, BEMs for combined field integral equations (CFIEs) on Lipschitz domains were investigated in [8, 25]. When the wave number belongs to the spectrum of the interior Maxwell’s problem, the EFIE and MFIE are not uniquely solvable, resulting in resonant instabilities of Galerkin discretization systems. CFIEs owe their name to an appropriate combination of the Maxwell single and double layer potentials, thereby getting rid of the spurious resonances. In [8], the authors exploited the coercivity of the EFIE operator by applying a compact regularization to the double layer part. The continuous and discrete variational problems were proven to be uniquely solvable at all wave numbers. A Galerkin boundary element discretization was proposed, whose numerical solutions satisfy an asymptotically quasi-optimal error estimate. The paper [25] introduced a preconditioned CFIE on which the EFIE and MFIE operators are preconditioned by their counterparts of a purely imaginary wave number. By means of a Calderón projection formula, the governing CFIE operator was cast into an equivalent formulation that, up to a compact perturbation, is $S_{i\kappa}$-coercive, with $S_{i\kappa}$ the EFIE operator of purely imaginary wave number $i\kappa$. A Galerkin discretization of the mixed variational problem using the lowest-order Raviart-Thomas space and its dual Buffa-Christiansen space was analyzed, producing quasi-optimal numerical solutions. Moreover, the resulting matrix systems were shown to be well-conditioned on fine meshes and stable at resonant frequencies. Nonetheless, the proposed discretization scheme in [25] is not the one typically used by practitioners. In practice, the MFIE part of the CFIE is commonly discretized using Raviart-Thomas basis functions for the solution space and Buffa-Christiansen functions for the test space [28, 27, 26].
@@
This paper aims to provide a rigorous analysis on BEMs for the MFIE on Lipschitz polyhedra. Like the EFIE and CFIEs, the MFIE is widely used in practical applications to model the electromagnetic scattering by perfect electric conductors. Despite its utility, however, attention paid to the MFIE is not at the same level with those for the EFIE and CFIEs. This is evident in the relatively limited number of papers devoting to BEMs for the MFIE, with most of them contributed by practitioners, see, e.g., [32, 18, 19, 15, 23, 24]. On smooth surfaces, the MFIE operator is well-known to be a Fredholm operator of second kind [7]. The existence and uniqueness of a solution to the continuous variational problem of the MFIE has been shown for smooth domains, provided that the wave number is not an eigenvalue of the interior Maxwell’s problem, see, e.g., [7]. In the discrete setting, this is not even necessarily true, since the boundary element subspace may not be dual to itself. More precisely, the Galerkin discretization of the identity operator (the principal part) occurring in the MFIE using lowest-order Raviart-Thomas elements has been proven to not satisfy the discrete inf-sup condition, see [11, Section 3.1]. As a result, the coercivity of the corresponding Galerkin discretization of the MFIE does not hold true.
@@
The contribution of this paper is twofold. Firstly, we prove that the continuous variational problem of the MFIE is also uniquely solvable on Lipschitz domains, of course provided that the wave number does not belong to the spectrum of the interior Maxwell’s problem. This is done by means of the elliptic operator $S_{i\kappa}$. More particularly, we show that the MFIE operator is $S_{i\kappa}$-coercive up to a compact perturbation. It is noteworthy that on non-smooth Lipschitz surfaces, the average of exterior and interior double layer boundary integral operators is not necessarily compact. Secondly, we analyze the mixed discretization scheme for the MFIE introduced in [15], when the boundary $\Gamma$ is Lipschitz continuous. The authors of [15] proposed a Galerkin discretization of the MFIE employing Raviart-Thomas basis functions for the solution space and Buffa-Christiansen functions for the test space, on which they exploited the duality between the Buffa-Christiansen space and the Raviart-Thomas space [2]. Given that the average double layer boundary integral operator is compact on smooth surfaces, this duality supplies the coercivity. Fortunately, the coercivity of the mixed discretization of the MFIE remains valid when the surface is solely Lipschitz continuous, thanks to the $S_{i\kappa}$-coercivity of the MFIE operator. As a consequence, the asymptotic quasi-optimality of numerical solutions is obtained. In addition, we show that the resulting matrix system is well-conditioned regardless of the boundary mesh refinement. To end this paragraph, we state here the following essential assumption and stop repeating it in the remainder of the paper. {assumption} The wave number is bounded away from the spectrum of the interior Maxwell’s problem.
@@
This paper is organized as follows. The next section recalls the standard functional setting for Maxwell’s equations on Lipschitz polyhedra, including the tangential trace operators, trace spaces, relevant tangential differential operators and an integration by parts formula. Section 3 gives the definition and properties of Maxwell potentials and boundary integral operators. The MFIE and its variational formulation are derived in Section 4. The $S_{i\kappa}$-coercivity of the MFIE operator is proven for Lipschitz domains, implying the unique solvability of the continuous variational problem. In Section 5, we analyze a Galerkin boundary element discretization of the MFIE. Its solvability, an asymptotically quasi-optimal error estimate for discrete solutions and the bounded condition number of the matrix system are established in this section. Some numerical results are presented in Section 6, aiming to verify the numerical analysis. Finally, we provide some conclusion remarks and discuss some possibilities for future works in Section 7.
@@ 2 Function spaces
Let $\Omega\subset\mathbb{R}^{3}$ be a bounded polyhedron with Lipschitz continuous boundary $\Gamma$, which only consists of flat faces $\Gamma_{j}$, $j=1,2,\ldots,N_{F}$, i.e., $\overline{\Gamma}=\bigcup_{j}\overline{\Gamma_{j}}$. The unbounded complement $\Omega^{c}:=\mathbb{R}^{3}\setminus\overline{\Omega}$ and the unit normal vector $\textbf{n}\in\operatorname{\mathbf{L}}^{\infty}(\Gamma)$ pointing from $\Omega$ to $\Omega^{c}$. On the polyhedral boundary $\Gamma$, the usual Sobolev spaces $\operatorname{H}^{s}(\Gamma)$ are defined invariantly for $s\in[0,1]$, with convention $\operatorname{H}^{0}(\Gamma)\equiv\operatorname{L}^{2}(\Gamma)$ [20, Section 1.3.3]. For $s>1$, the spaces $\operatorname{H}^{s}(\Gamma)$ can be characterized face by face, with a careful treatment at their common edges, see [3, 4]. We denote by $\operatorname{H}^{-s}(\Gamma)$ the dual space of $\operatorname{H}^{s}(\Gamma)$, and $\left\langle{\cdot,\cdot}\right\rangle_{s,\Gamma}$ their duality pairing. In addition, we introduce the space $\operatorname{H}^{3/2}_{\sim}(\Gamma):=\gamma(\operatorname{H}^{2}(\Omega))$ and its dual $\operatorname{H}^{-3/2}_{\sim}(\Gamma)$, with $\gamma$ the standard trace operator. The duality pairing between $\operatorname{H}_{\sim}^{-3/2}(\Gamma)$ and $\operatorname{H}_{\sim}^{3/2}(\Gamma)$ is denoted by $\left\langle{\cdot,\cdot}\right\rangle_{\sim,3/2,\Gamma}$. Here, we make the usual identification $\operatorname{L}^{2}(\Gamma)\subset\operatorname{H}^{-s}(\Gamma)$ and $\operatorname{L}^{2}(\Gamma)\subset\operatorname{H}^{-3/2}_{\sim}(\Gamma)$ via the $\operatorname{L}^{2}(\Gamma)$-pairing.
@@
$(\gamma_{D}\boldsymbol{u})(\boldsymbol{x})=\boldsymbol{u}(\boldsymbol{x})\times\textbf{n}(\boldsymbol{x}),$
for almost all $\boldsymbol{x}\in\Gamma$. This operator can be extended to a continuous mapping from $\operatorname{\mathbf{H}}(\textbf{{curl}},\Omega)$ to $\operatorname{\mathbf{H}}^{-1/2}(\Gamma)$, see [20, Theorem 1.5.1.1]. In order to determine the actual range of $\gamma_{D}$, we adopt the definition $\operatorname{\mathbf{H}}^{s}_{\times}(\Gamma):=\gamma_{D}(\operatorname{\mathbf{H}}^{s+1/2}(\Omega))$, with $s\in(0,1)$, from [8, Definition 2.1]. For $s=0$, we set the space $\operatorname{\mathbf{H}}^{0}_{\times}(\Gamma)$ identical with
@@
$\left\langle{\boldsymbol{u},\boldsymbol{v}}\right\rangle_{\times,\Gamma}:=\int_{\Gamma}(\boldsymbol{u}\times\textbf{n})\cdot\boldsymbol{v}\,\mathrm{d}s,$
for all $\boldsymbol{u},\boldsymbol{v}\in\operatorname{\mathbf{L}}^{2}_{\textbf{t}}(\Gamma)$. The dual space of $\operatorname{\mathbf{H}}^{s}_{\times}(\Gamma)$ is denoted by $\operatorname{\mathbf{H}}^{-s}_{\times}(\Gamma)$, whose elements are identified via the $\operatorname{\mathbf{L}}^{2}_{\textbf{t}}(\Gamma)$-pairing ,×$\left\langle{\cdot,\cdot}\right\rangle_{\times,\Gamma}$.
Let us introduce relevant surface differential operators. The surface operator $\textbf{{curl}}_{\Gamma}:\widetilde{\operatorname{H}}^{3/2}(\Gamma)\to\operatorname{\mathbf{H}}^{1/2}_{\times}(\Gamma)$ can be defined by
$\textbf{{curl}}_{\Gamma}\gamma(\varphi)=\gamma_{D}(\textbf{grad}\varphi),$
for all $\varphi\in\operatorname{H}^{2}(\Omega)$, see [6, Proposition 3.4]. Then, the surface divergence operator $\textup{div}_{\Gamma}:\operatorname{\mathbf{H}}^{-1/2}_{\times}(\Gamma)\to\operatorname{H}^{3/2}_{\sim}(\Gamma)$ is defined as dual to $\textbf{{curl}}_{\Gamma}$ in the sense
$\left\langle{\textup{div}_{\Gamma}\boldsymbol{u},\varphi}\right\rangle_{\sim,3/2,\Gamma}=-\left\langle{\boldsymbol{u},\textbf{{curl}}_{\Gamma}\varphi}\right\rangle_{1/2,\Gamma},$
for all $\boldsymbol{u}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\Gamma)$ and $\varphi\in\widetilde{\operatorname{H}}^{3/2}(\Gamma)$. Now, we are able to introduce the space
@@
The next theorem presents the extension of the $\operatorname{\mathbf{L}}^{2}_{\textbf{t}}(\Gamma)$-pairing ,×$\langle{\cdot,\cdot}\rangle_{\times,\Gamma}$ to the space $\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$ and its self-duality with respect to this pairing, see, e.g., [4] and [6, Lemma 5.6].
@@
The tangential trace operator $\gamma_{D}$ maps from $\operatorname{\mathbf{H}}(\textbf{{curl}},\Omega)$ onto $\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$, making it the natural trace space for electromagnetic fields. The following theorem provides an integration by parts formula, which is essential for further analysis in this paper, see [6, Theorem 4.1] or [13, Theorem 2.3].
@@
The tangential trace operator $\gamma_{D}$ continuously maps from $\operatorname{\mathbf{H}}(\textbf{{curl}},\Omega)$ onto $\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$, which possesses a continuous right inverse. In addition, the following integration by parts formula holds for all $\boldsymbol{u},\boldsymbol{v}\in\operatorname{\mathbf{H}}(\textbf{{curl}},\Omega)$
(1) $\int_{\Omega}(\textbf{{curl}}\,\boldsymbol{u}\cdot\boldsymbol{v}-\boldsymbol{u}\cdot\textbf{{curl}}\,\boldsymbol{v})\,\mathrm{d}\boldsymbol{x}=-\left\langle{\gamma_{D}\boldsymbol{u},\gamma_{D}\boldsymbol{v}}\right\rangle_{\times,\Gamma}.$

### v2

@@ Abstract
This paper provides a rigorous analysis of boundary element methods for the magnetic field integral equation on Lipschitz polyhedra. The magnetic field integral equation is widely used in practical applications to model electromagnetic scattering by a perfectly conducting body. The governing operator is shown to be coercive by means of the electric field integral operator with a purely imaginary wave number. Consequently, the continuous variational problem is uniquely solvable, given that the wave number does not belong to the spectrum of the interior Maxwell’s problem. A Petrov-Galerkin discretization scheme is then introduced, employing Raviart-Thomas boundary elements for the solution space and Buffa-Christiansen boundary elements for the test space. Under a mild assumption depending only on the geometrical domain, the corresponding discrete inf-sup condition is proven, implying the unique solvability of the discrete problem. An asymptotically quasi-optimal error estimate for numerical solutions is established, and the convergence rate of the numerical scheme is examined. In addition, the resulting matrix system is shown to be well-conditioned regardless of the mesh refinement. Some numerical results are presented to support the theoretical analysis.
@@
magnetic field integral equation, MFIE, electromagnetic scattering, boundary element methods, Petrov-Galerkin boundary element discretization
@@ 1 Introduction
With boundary element methods (BEMs) in mind, we assume that the domain $\Omega$ is a polyhedron with Lipschitz continuous boundary $\Gamma:=\partial\Omega$ that only consists of flat faces $\Gamma_{j}$, $j=1,2,\ldots,N_{F}$. The extension of the results to Lipschitz curvilinear polyhedra is straightforward. On polyhedral boundary $\Gamma$, the definition of Sobolev spaces $\operatorname{H}^{s}(\Gamma)$, in particular with $\left|{s}\right|\geq 1$, and differential operators are not obvious as those on smooth surfaces. By localization to each face $\Gamma_{j}$, Buffa and Ciarlet successfully developed functional setting for Maxwell’s equations on Lipschitz polyhedra, for which tangential traces of vector fields in $\operatorname{\mathbf{H}}(\textbf{{curl}},\Omega)$, their trace spaces and relevant surface differential operators were appropriately characterized, see [2, 4, 5]. All ingredients necessary to develop a standard functional theory, including integration by parts formulas and Hodge decompositions, were also provided. A generalization for general Lipschitz domains was then presented in [7].
@@
Based on the developed functional theory for Maxwell’s equations on Lipschitz domains, a natural BEM for the EFIE on polyhedra was studied in [26]. The continuous and discrete variational problems are coercive with respect to continuous and semi-discrete Hodge decompositions, respectively. The Galerkin discretization of the EFIE employing Raviart-Thomas boundary elements was analyzed, exhibiting the quasi-optimal convergence of numerical solutions. Another discretization scheme for the EFIE on non-smooth Lipschitz domains was proposed in [6], using standard boundary elements rather than Raviart-Thomas elements. On polyhedral surfaces, an asymptotically quasi-optimal convergence behaviour was also achieved. BEMs for Maxwell transmission problems on Lipschitz domains were investigated in [10]. The Galerkin discretization with Raviart-Thomas spaces was considered, and the results were applied to electromagnetic scattering by perfect electric conductors and by dielectric bodies. In the case of the perfect conductor scattering, only the EFIE was examined, also yielding quasi-optimal numerical solutions. To determine the actual rate of convergence on polyhedral domains, the authors of [6] and [26] all took into account the singularities of solution to Maxwell’s equations, which were investigated in [19]. More specifically, the convergence rate of BEMs for the EFIE is limited by the regularity parameter $s^{\ast}$ of the Laplace-Beltrami operator on $\Gamma$, which depends on the geometry of $\Gamma$ in neighborhoods of vertices only.
@@
In addition to the EFIE, BEMs for combined field integral equations (CFIEs) on Lipschitz domains were investigated in [9, 29]. When the wave number belongs to the spectrum of the interior Maxwell’s problem, the EFIE and MFIE are not uniquely solvable, resulting in resonant instabilities of Galerkin discretization systems. CFIEs owe their name to an appropriate combination of the Maxwell single and double layer potentials, thereby getting rid of the spurious resonances. In [9], the authors exploited the coercivity of the EFIE operator by applying a compact regularization to the double layer part. The continuous and discrete variational problems were proven to be uniquely solvable at all wave numbers. A Galerkin boundary element discretization was proposed, whose numerical solutions satisfy an asymptotically quasi-optimal error estimate. The paper [29] introduced a preconditioned CFIE on which the EFIE and MFIE operators are preconditioned by their counterparts of a purely imaginary wave number. By means of a Calderón projection formula, the governing CFIE operator was cast into an equivalent formulation that, up to a compact perturbation, is $S_{i\kappa}$-coercive, with $S_{i\kappa}$ the EFIE operator of purely imaginary wave number $i\kappa$. A Galerkin discretization of the mixed variational problem using the lowest-order Raviart-Thomas space and its dual Buffa-Christiansen space was analyzed, producing quasi-optimal numerical solutions. Moreover, the resulting matrix systems were shown to be well-conditioned on fine meshes and stable at resonant frequencies. Nonetheless, the proposed discretization scheme in [29] is not the one typically used by practitioners. In practice, the MFIE part of CFIEs is commonly discretized using Raviart-Thomas boundary elements for the solution space and Buffa-Christiansen elements for the test space. For some examples, the reader is referred to [1, 32, 30, 31].
@@
This paper aims to provide a rigorous analysis of BEMs for the MFIE on Lipschitz polyhedra. Like the EFIE and CFIEs, the MFIE is widely used in practical applications to model the electromagnetic scattering by perfect electric conductors. Despite its utility, however, attention paid to the MFIE is not at the same level with those for the EFIE and CFIEs. This is evident in the relatively limited number of papers devoting to BEMs for the MFIE, with most of them contributed by practitioners, see, e.g., [36, 20, 21, 17, 27, 28]. On smooth surfaces, the MFIE operator is well-known to be a Fredholm operator of second kind [8]. The existence and uniqueness of a solution to the continuous variational problem of the MFIE has been shown for smooth domains, provided that the wave number is not an eigenvalue of the interior Maxwell’s problem, see, e.g., [8]. In the discrete setting, the existence does not necessarily hold true, since the boundary element subspace may not be dual to itself. More precisely, the Galerkin discretization of the identity operator (the principal part of the MFIE operator) using lowest-order Raviart-Thomas elements has been proven to not satisfy the discrete inf-sup condition, see [13, Section 3.1]. As a result, the coercivity of the corresponding Galerkin discretization of the MFIE does not hold true.
@@
The contribution of this paper is twofold. First, we prove that the continuous variational problem of the MFIE is also uniquely solvable on Lipschitz domains, of course provided that the wave number does not belong to the spectrum of the interior Maxwell’s problem. More particularly, we show that the MFIE operator is $S_{i\kappa^{\prime}}$-coercive up to a compact perturbation, for any $\kappa^{\prime}>0$. It is noteworthy that on non-smooth Lipschitz surfaces, the average of exterior and interior traces of the double layer potential operator is not necessarily compact. Secondly, we analyze the mixed discretization scheme for the MFIE introduced in [17], when the boundary $\Gamma$ is Lipschitz continuous. The authors of [17] proposed a Petrov-Galerkin discretization of the MFIE employing Raviart-Thomas boundary elements for the solution space and Buffa-Christiansen boundary elements for the test space, on which they exploited the duality between the Buffa-Christiansen space and the Raviart-Thomas space [3]. Given that the average double layer boundary integral operator is compact on smooth surfaces, this duality supplies the coercivity. When the surface $\Gamma$ is solely Lipschitz continuous, the coercivity of the mixed discretization remains valid if an additional stability condition is satisfied, which only depends on the boundary $\Gamma$. This condition is mild as it is fulfilled for typical geometries in practical applications. As a consequence, the asymptotic quasi-optimality of numerical solutions is obtained. In addition, we show that the resulting matrix system is well-conditioned regardless of the boundary mesh refinement. To end this paragraph, we state here the essential assumption on the wave number of Maxwell’s equations and stop repeating it in the remainder of the paper. {assumption} The wave number is bounded away from the spectrum of the interior Maxwell’s problem.
@@
This paper is organized as follows. The next section recalls the standard functional setting for Maxwell’s equations on Lipschitz polyhedra, including the tangential trace operators, trace spaces, relevant tangential differential operators and an integration by parts formula. Section 3 gives the definition and properties of Maxwell potentials and boundary integral operators. The MFIE and its appropriate variational formulation are derived in Section 4. The governing MFIE operator is shown to be compact perturbation of a $S_{i\kappa^{\prime}}$-coercive operator, implying the unique solvability of the continuous variational problem. In Section 5, we analyze a Galerkin boundary element discretization of the MFIE. Its solvability under a stability assumption, an asymptotically quasi-optimal error estimate for discrete solutions, and the bounded condition number of the matrix system are established in this section. The numerical verification of the stability condition is discussed in Section 6. Section 7 presents some numerical results aiming to support the numerical analysis. Finally, we provide some conclusion remarks and discuss some possibilities for future works in Section 8.
@@ 2 Function spaces
Let $\Omega\subset\mathbb{R}^{3}$ be a bounded polyhedron with Lipschitz continuous boundary $\Gamma$, which only consists of flat faces $\Gamma_{j}$, $j=1,2,\ldots,N_{F}$, i.e., $\overline{\Gamma}=\bigcup_{j}\overline{\Gamma_{j}}$. The unbounded complement $\Omega^{c}:=\mathbb{R}^{3}\setminus\overline{\Omega}$ and the unit normal vector $\textbf{n}\in\operatorname{\mathbf{L}}^{\infty}(\Gamma)$ pointing from $\Omega$ to $\Omega^{c}$. On the polyhedral boundary $\Gamma$, the usual Sobolev spaces $\operatorname{H}^{s}(\Gamma)$ are defined invariantly for $s\in[0,1]$, with convention $\operatorname{H}^{0}(\Gamma)\equiv\operatorname{L}^{2}(\Gamma)$ [23, Section 1.3.3]. For $s>1$, the spaces $\operatorname{H}^{s}(\Gamma)$ can be characterized face by face, with a careful treatment at their common edges, see [4, 5]. We denote by $\operatorname{H}^{-s}(\Gamma)$ the dual space of $\operatorname{H}^{s}(\Gamma)$, and $\left\langle{\cdot,\cdot}\right\rangle_{s,\Gamma}$ their duality pairing. In addition, we introduce the space $\widetilde{\operatorname{H}}^{3/2}(\Gamma):=\gamma(\operatorname{H}^{2}(\Omega))$ and its dual $\widetilde{\operatorname{H}}^{-3/2}(\Gamma)$, with $\gamma$ the standard trace operator [^1] The tilde is used to distinguish the trace of $\operatorname{H}^{2}(\Omega)$ on the boundary $\Gamma$ and the space $\operatorname{H}^{3/2}(\Gamma)$, as they are not identical in the case of a non-smooth Lipschitz domain $\Omega$.. The duality pairing between $\widetilde{\operatorname{H}}^{-3/2}(\Gamma)$ and $\widetilde{\operatorname{H}}^{3/2}(\Gamma)$ is denoted by $\left\langle{\cdot,\cdot}\right\rangle_{\sim,3/2,\Gamma}$. Here, we make the usual identification $\operatorname{L}^{2}(\Gamma)\subset\operatorname{H}^{-s}(\Gamma)$ and $\operatorname{L}^{2}(\Gamma)\subset\widetilde{\operatorname{H}}^{-3/2}(\Gamma)$ via the $\operatorname{L}^{2}(\Gamma)$-pairing.
@@
$(\gamma_{D}\boldsymbol{u})(\boldsymbol{x}):=\boldsymbol{u}(\boldsymbol{x})\times\textbf{n}(\boldsymbol{x}),$
for almost all $\boldsymbol{x}\in\Gamma$. This operator can be extended to a continuous mapping from $\operatorname{\mathbf{H}}(\textbf{{curl}},\Omega)$ to $\operatorname{\mathbf{H}}^{-1/2}(\Gamma)$, see [23, Theorem 1.5.1.1]. In order to determine the actual range of $\gamma_{D}$, we adopt the definition $\operatorname{\mathbf{H}}^{s}_{\times}(\Gamma):=\gamma_{D}(\operatorname{\mathbf{H}}^{s+1/2}(\Omega))$, with $s\in(0,1)$, from [9, Definition 2.1]. For $s=0$, we set the space $\operatorname{\mathbf{H}}^{0}_{\times}(\Gamma)$ identical with
@@
$\left\langle{\boldsymbol{u},\boldsymbol{v}}\right\rangle_{\times,\Gamma}:=\int_{\Gamma}(\boldsymbol{u}\times\textbf{n})\cdot\boldsymbol{v}\,\mathrm{d}s,\qquad\qquad\forall\boldsymbol{u},\boldsymbol{v}\in\operatorname{\mathbf{L}}^{2}_{\textbf{t}}(\Gamma).$
The dual space of $\operatorname{\mathbf{H}}^{s}_{\times}(\Gamma)$ is denoted by $\operatorname{\mathbf{H}}^{-s}_{\times}(\Gamma)$, whose elements are identified via the $\operatorname{\mathbf{L}}^{2}_{\textbf{t}}(\Gamma)$-pairing ,×$\left\langle{\cdot,\cdot}\right\rangle_{\times,\Gamma}$.
Let us introduce relevant surface differential operators. The surface operator $\textbf{{curl}}_{\Gamma}:\widetilde{\operatorname{H}}^{3/2}(\Gamma)\to\operatorname{\mathbf{H}}^{1/2}_{\times}(\Gamma)$ can be defined by [7, Proposition 3.4]
$\textbf{{curl}}_{\Gamma}\gamma(\varphi)=\gamma_{D}(\textbf{grad}\varphi),\qquad\qquad\quad\forall\varphi\in\operatorname{H}^{2}(\Omega).$
Then, the surface divergence operator $\textup{div}_{\Gamma}:\operatorname{\mathbf{H}}^{-1/2}_{\times}(\Gamma)\to\widetilde{\operatorname{H}}^{-3/2}(\Gamma)$ is defined as dual to $\textbf{{curl}}_{\Gamma}$ in the sense
$\left\langle{\textup{div}_{\Gamma}\boldsymbol{u},\varphi}\right\rangle_{\sim,3/2,\Gamma}=-\left\langle{\boldsymbol{u},\textbf{{curl}}_{\Gamma}\varphi}\right\rangle_{1/2,\Gamma},\qquad\quad\forall\boldsymbol{u}\in\operatorname{\mathbf{H}}^{-1/2}_{\times}(\Gamma),\varphi\in\widetilde{\operatorname{H}}^{3/2}(\Gamma).$
Now, we are able to introduce the space
@@
The next theorem presents the extension of the $\operatorname{\mathbf{L}}^{2}_{\textbf{t}}(\Gamma)$-pairing ,×$\langle{\cdot,\cdot}\rangle_{\times,\Gamma}$ to a bilinear form on the space $\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$ and its self-duality with respect to this pairing, see, e.g., [5] and [7, Lemma 5.6].
@@
The tangential trace operator $\gamma_{D}$ maps from $\operatorname{\mathbf{H}}(\textbf{{curl}},\Omega)$ onto $\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$, making $\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$ the natural trace space for electromagnetic fields. The following theorem provides an integration by parts formula for functions in $\operatorname{\mathbf{H}}(\textbf{{curl}},\Omega)$, which is taken from [7, Theorem 4.1] and [15, Theorem 2.3].
@@
The tangential trace operator $\gamma_{D}$ continuously maps from $\operatorname{\mathbf{H}}(\textbf{{curl}},\Omega)$ onto $\operatorname{\mathbf{H}}^{-1/2}_{\times}(\textup{div}_{\Gamma},\Gamma)$, which possesses a continuous right inverse. In addition, the following integration by parts formula holds
(1) $\int_{\Omega}(\textbf{{curl}}\,\boldsymbol{u}\cdot\boldsymbol{v}-\boldsymbol{u}\cdot\textbf{{curl}}\,\boldsymbol{v})\,\mathrm{d}\boldsymbol{x}=-\left\langle{\gamma_{D}\boldsymbol{u},\gamma_{D}\boldsymbol{v}}\right\rangle_{\times,\Gamma},\qquad\forall\boldsymbol{u},\boldsymbol{v}\in\operatorname{\mathbf{H}}(\textbf{{curl}},\Omega).$